Find the area of triangle ABC if A=56 degrees, b=20 ft, and c=12 feet. Round to the nearest tenth.
Answer by Alan3354(21605)

Find the area of triangle ABC if A=56 degrees, b=20 ft, and c=12 feet. Round to the nearest tenth.
Draw AC as the base (horizontal), which is side b = 20. Draw the altitude from vertex B to the base, call it point D. That makes a right triangle ABD, and BD = 12*sin(56).
The area is bh/2, = AC*BD/2
=20*12*sin(56)/2
=120*sin(56)
=99.5 square units